OpenStudy (anonymous):

please help! i only have 4 days of school left!!!

OpenStudy (anonymous):

Find the length of AB and the midpoint of AB. Round the length of AB to the nearest tenth if necessary. 1.A(0,-2) B(-6,-9) 2.A(8,11) B(-5,2)

OpenStudy (jim766):

to find length \[\sqrt{(x-x)^2+(y-y)^2}\] do you know how to use it?

OpenStudy (jim766):

\[\sqrt{(0--6)^2+(-2--9)^2}\] \[\sqrt{6^2+5^2} = \sqrt{36+25}=\sqrt{61}=7.8\]

OpenStudy (jim766):

to find mid point add up the x's of the points add up the ys of the points divide each by 2 A (0 , -2) B (-6, -9) ------ add them -6, -11 now divide each by 2 -- -- 2 2 (-3, -5.5) mid point

OpenStudy (jim766):

does that make sense? do #2 the same way

OpenStudy (anonymous):

yes thank you so much